OpenStudy (anonymous):

PLEASEEEEE HELP!!!!! The volume of a cone is 183.26 cubic feet and its radius is 5 feet. What is the cone's height? Round to the nearest foot.

OpenStudy (anonymous):

volume of a cone V=1/3 pi r^2 h

OpenStudy (anonymous):

so its 183.26=1/3(3.14)5^2

OpenStudy (anonymous):

it should be 183.86=1/3(3.14) 5^2 h then calculate h

OpenStudy (anonymous):

thats what i had ? and do i multiply (3.14) and 1/3 ??

OpenStudy (anonymous):

183.26*3=3.14*25*h

OpenStudy (anonymous):

549.78 for the left side and other is 78.5 ??

OpenStudy (anonymous):

h=549.78/78.5

OpenStudy (anonymous):

7.003566879

OpenStudy (anonymous):

approx. 7

OpenStudy (anonymous):

how do you do this one ?? Marissa has a snow cone cart that serves about 30 people each hour at the local beach. If the cone shaped snow cone cups have a height of 3 inches and a radius of 1.5 inches, what volume of shaved ice is needed each hour? Round to the nearest cubic inch.